# Approximate predictive control barrier function for discrete-time systems

### v1

@@ 3 PREDICTIVE CONTROL BARRIER FUNCTIONS
Given the current state $x(k)$, a trajectory of states $x_{i|k}$ and inputs $u_{i|k}$ is planned according to the system dynamics in (9c) such that the state and input constraints (2) are satisfied for all predicted time steps. We note that as the initial state can lie outside the state constraints $\mathcal{X}$, the state constraints are formulated in terms of soft constraints with respect to slack variables $\xi^{*}_{i|k}$ which are computed according to a second optimisation problem discussed later in this section. The state constraints are subject to an additional tightening, i.e., $\overline{\mathcal{X}}_{i}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq-\Delta_{i}+\xi_{i|k}\}$ with $\Delta_{i}-\Delta_{i-1}>\Delta_{0}=0$, in order to establish stability guarantees of the closed-loop. A terminal constraint (9f) is imposed on the final predicted state, where $h_{f}$ is a CBF according to Definition 1. As the explicit formulation of $h_{f}$ is used in the optimisation problem, it is assumed to be known.
@@
A terminal control barrier function $h_{f}:\mathcal{D}_{f}\rightarrow\mathbb{R}$ with domain $\mathcal{D}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq\gamma_{f}\}$ and corresponding safe set $\mathcal{S}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq 0\}\subseteq\overline{\mathcal{X}}_{N-1}(0)$ exists according to Definition 1 for the system dynamics (1) and is known.
@@
Building upon the introduced predictive safety filter formulation, it is shown in [12], that a similar enhancement mechanism can be formulated for CBFs to increase the domain of attraction of the closed-loop system. As in practical applications, the initial state $x(0)$ may not be feasible without soft constraints or unmodelled external disturbances could result in infeasible states for the optimisation problem, a recovery mechanism is required, which ensures that the system converges back to an invariant subset of the constraints $\mathcal{X}$ in such events. Through their property of stability of invariant sets in a given domain, CBFs offer a theoretical framework which is suitable for such a recovery mechanism. This enhancement property for the domain of attraction of the closed-loop system is achieved by computing the slack variables $\xi^{*}_{i|k}$ in (9) according to the following optimisation problem, which minimises the state constraint violations and a weighted penalty on the terminal slack $\xi_{N|k}$, i.e.
$h_{\mathrm{PB}}(x(k)):=\min_{u_{i|k},\xi_{i|k}}\;\;$ $\alpha_{f}\xi_{N|k}+\sum_{i=0}^{N-1}\lvert\lvert{\xi_{i|k}}\rvert\rvert$ (10a)
@@
[…]
The optimal value function of (10), i.e. $h_{\mathrm{PB}}(x)$, is then shown to be a CBF according to Definition 1 in [12].
@@
Invariance of the safe set $\mathcal{S}_{\mathrm{PB}}$ in Theorem 2 results from recursive feasibility of the problem with $\xi^{*}_{i|k}=0$ from standard arguments, while the existence of a continuous, positive decrease $\Delta h_{\mathrm{PB}}$ in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ is ensured through the terminal CBF $h_{f}$ in Assumption 2 if $\xi^{*}_{N|k}>0$ and through the additional tightening $\Delta_{i}$ of the state constraints $\overline{\mathcal{X}}_{i}(\xi_{i|k})$ if $\xi^{*}_{N|k}=0$. Clearly, if $\xi^{*}_{i|k}=0$, the problem (9) is solved with additional tightening on the state constraints and the corresponding tightenings $\Delta_{i}$ should therefore be chosen small.

### v2

@@ 3 PRELIMINARIES ON PREDICTIVE CONTROL BARRIER FUNCTIONS
Given the current state $x(k)$, a nominal trajectory of states $x_{i|k}$ and inputs $u_{i|k}$ is planned according to the system dynamics in (8c) such that the state and input constraints (8d) and (8e) are satisfied for all predicted time steps. We note that as the state can lie outside the set $\mathcal{X}$, the state constraints are formulated in terms of soft constraints with respect to slack variables $\xi^{*}_{i|k}$ which are computed according to a second optimisation problem discussed later in this section. The state constraints are subject to an additional tightening, i.e., $\overline{\mathcal{X}}_{i}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq-\Delta_{i}+\xi_{i|k}\}$ with $\Delta_{i}>\Delta_{i-1}>\Delta_{0}=0$, in order to establish stability guarantees of the closed-loop. A terminal constraint (8f) is imposed on the final predicted state, where $h_{f}:\mathcal{D}_{f}\rightarrow\mathbb{R}$ is a CBF according to Definition 1 and is assumed to be known.
@@
A terminal CBF $h_{f}:\mathcal{D}_{f}\rightarrow\mathbb{R}$ with domain $\mathcal{D}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq\gamma_{f}\}$ for $\gamma_{f}>0$ and corresponding safe set $\mathcal{S}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq 0\}\subseteq\overline{\mathcal{X}}_{N-1}(0)$ exists according to Definition 1 for the dynamics (1) and is known.
@@
Building upon the introduced predictive safety filter formulation, it is shown in [12], that a similar enhancement mechanism can be formulated for CBFs to increase the domain of attraction of the closed-loop system. As in practical applications, the initial state $x(0)$ may not be feasible without soft constraints or unmodelled external disturbances could result in infeasible states for the optimisation problem, a recovery mechanism is required, which ensures that the system converges back to an invariant subset of the constraints $\mathcal{X}$ in such events. Through their property of stability of invariant sets in a given domain, CBFs offer a theoretical framework which is suitable for such a recovery mechanism. Such an enhancement property for the domain of attraction of the closed-loop system is achieved by computing the slack variables $\xi^{*}_{i|k}$ in (8) according to the following optimisation problem, which minimises the state constraint violations and a weighted penalty on the terminal slack $\xi_{N|k}$ with $\alpha_{f}>0$, i.e.
$\hskip-2.84544pth_{\mathrm{PB}}(x(k)):=\min_{x_{i|k},u_{i|k},\xi_{i|k}}\;\;$ $\alpha_{f}\xi_{N|k}+\sum_{i=0}^{N-1}\lvert\lvert{\xi_{i|k}}\rvert\rvert$ (9a)
@@
$x_{i|k}\in\overline{\mathcal{X}}_{i}(\xi_{i|k}),\;0\leq\xi_{i|k},$ (9e)
$h_{f}(x_{N|k})\leq\xi_{N|k},\;0\leq\xi_{N|k}.$ (9f)
In [12], it is shown that the optimal value function of (9), i.e., $h_{\mathrm{PB}}(x)$, is a CBF according to Definition 1 if $\alpha_{f}$ is chosen sufficiently large.
@@
Invariance of the safe set $\mathcal{S}_{\mathrm{PB}}$ in Theorem 2 results from recursive feasibility of the problem with $\xi^{*}_{i|k}=0$ from standard arguments, while the existence of a continuous, positive decrease $\Delta h_{\mathrm{PB}}$ in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ is ensured through the terminal CBF $h_{f}$ in Assumption 2 if $\xi^{*}_{N|k}>0$ and through the additional tightening $\Delta_{i}$ of the state constraints $\overline{\mathcal{X}}_{i}(\xi_{i|k})$ if $\xi^{*}_{N|k}=0$. Clearly, if $\xi^{*}_{i|k}=0$, the problem (8) is solved with additional tightening on the state constraints and the corresponding tightenings $\Delta_{i}$ should therefore be chosen small. An illustration of the different set definitions can be found in [12, Figure 4].
4 EXPLICIT APPROXIMATION OF THE PCBF
In this section, we present the main proposed approximation-based algorithm. Furthermore, we propose a continuous extension for the PCBF, which is guaranteed to be negative for states in the safe set $\mathcal{S}_{\mathrm{PB}}$, i.e., when $h_{\mathrm{PB}}(x)=0$, thereby reducing safety filter interventions caused by approximation errors within $\mathcal{S}_{\mathrm{PB}}$.
4.1 Approximate PCBF
The optimal value function in (9) is a CBF according to Definition 4 and encodes the information of the entire prediction horizon of the predictive optimisation problem with respect to potential constraint violations. However, an explicit formulation of $h_{\mathrm{PB}}$ is generally not available. We therefore propose approximating the PCBF value function with a continuous regressor $\hat{h}_{\mathrm{PB}}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that
$\hat{h}_{\mathrm{PB}}(x)\approx h_{\mathrm{PB}}(x)\;\forall x\in\mathcal{D}_{\mathrm{PB}}.$ (10)
Due to the continuity property of the PCBF in (9) and the fact that it can be sampled offline, the function is suitable for approximation, such that a computationally efficient algorithm can be deployed. As the approximation $\hat{h}_{\mathrm{PB}}$ will be embedded in an optimisation problem, we consider neural networks in this paper due to the parametric nature of the regressor, allowing for a formulation independent of the number of training samples, compared to, e.g., kernel-based methods. Note that while in principle a safety filter policy $\pi(x,u_{\mathrm{p}})$ could be learned, this policy is generally not only non-continuous and multi-valued, but also has a higher input dimensionality for the data collection and training as $\pi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. As an illustrative example, consider a car approaching an obstacle. If the performance input is given to steer straight towards the obstacle, an optimal solution for the safety filter problem is given by either steering fully to the left or to the right.
In order to train the regressor for the PCBF approximation, a dataset of samples has to be collected offline in the form $\mathbb{D}=\{(x_{i},h_{\mathrm{PB}}(x_{i}))\}_{i=0}^{n_{d}}$. This data set can be sampled from the entire region of attraction $\mathcal{D}_{\mathrm{PB}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}\}$, where $\gamma_{f}$ is the sublevel set of the domain of the terminal CBF $\mathcal{D}_{f}$ and $\alpha_{f}$ is the terminal penalty in (9a), or from a subset deemed relevant to the specific application. In the Appendix, we propose a geometric sampling method based on [41], which allows efficiently sampling the PCBF by performing a pseudo random walk over a preselected domain, using previous solutions as warmstarts to the next computed optimisation problem.
Given a trained explicit approximation $\hat{h}_{\mathrm{PB}}$ of $h_{\mathrm{PB}}$, the corresponding CBF decrease condition (3a) can directly be used as a constraint similar to CBF-based safety filters in [16]. The optimisation problem which is solved at every time step is then given by
$\min_{u}$ $\|u-u_{\mathrm{p}}(k)\|$ (11)
s.t. $\hat{h}_{\mathrm{PB}}(f(x(k),u,0))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k)),$
$u\in\mathcal{U},$
where $\Delta\hat{h}_{\mathrm{PB}}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a decrease function of the approximate PCBF $\hat{h}_{\mathrm{PB}}(x)$ and $c_{\Delta}$ is a user-defined hyperparameter which is 1 if $\Delta\hat{h}_{\mathrm{PB}}(x)<0$ and can be chosen as $c_{\Delta}\in(0,1]$ otherwise. The main advantage of the proposed optimisation problem compared to the PCBF algorithm in [12], which solves (9) and (8) consecutively, is that the formulation in (11) is independent of the prediction horizon of the predictive control-based optimisation problems. Instead, only a single decrease constraint needs to be satisfied along with the input constraints and the only optimisation variable is $u\in\mathbb{R}^{m}$.
Given that the decrease function of $h_{\mathrm{PB}}(x)$, i.e., $\Delta h_{\mathrm{PB}}(x)$, is not explicitly available in a similar fashion to $h_{\mathrm{PB}}(x)$ itself, we propose an optimisation-based formulation for $\Delta\hat{h}_{\mathrm{PB}}$. The approximate decrease which is used in (11) is defined as
$\Delta\hat{h}_{\mathrm{PB}}(x):=\max_{u\in\mathcal{U}}\left(\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u,0))\right)$ (12)
From this formulation of the approximate decrease function, it follows that choosing the hyperparameter $c_{\Delta}=1$, requires the approximate PCBF $\hat{h}_{\mathrm{PB}}$ to undergo the maximal possible decrease. By selecting $c_{\Delta}\in(0,1)$ when a decrease can be achieved, safety interventions of the proposed filter can be reduced by relaxing the amount that the approximate PCBF is required to decrease. This hyperparameter can be tuned after obtaining $\hat{h}_{\mathrm{PB}}$ for desired closed-loop behaviour and provides an additional degree of freedom compared to the the algorithms in [12] and [13], where at every time step the maximal decrease is required to be achieved.
Remark 2
An additional advantage of the explicit approximation of the PCBF is given by the fact that proposed inputs $u_{\mathrm{p}}(k)$ can be very efficiently certified in terms of satisfying optimality with respect to (11) if $u_{\mathrm{p}}(k)\in\mathcal{U}$. It suffices to perform a simple evaluation, whether $\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k),0))\leq 0$, as it then holds that $u_{\mathrm{p}}(k)$ is an optimal solution to (11). This function evaluation can thereby reduce the need for solving two optimisation problems online at every time step. Additionally, it can avoid local minima when solving optimisation problems online, especially compared to predictive optimisation problems. The main proposed algorithm is then given in Algorithm 1, where first, the proposed input is certified and if the input is not deemed to be safe, the maximal decrease in (12) is computed and then used in the safety filter (11).
Algorithm 1 Approximate PCBF safety filter
1:for $k\geq 0$ do
2:  Measure state $x(k)$
3:  Obtain the nominal control input $u_{\mathrm{p}}(k)$
4:  if $u_{\mathrm{p}}(k)\in\mathcal{U}\wedge\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k)),0)\leq 0$ then
5:   Apply $u(k)\leftarrow u_{\mathrm{p}}(k)$ // nominal input is safe
6:  else
7:   Compute $\Delta\hat{h}_{\mathrm{PB}}(x(k))$ using (12)
8:   Solve (11) and obtain a solution $u^{*}$
9:   Apply $u(k)\leftarrow u^{*}$ // apply filtered input
10:  end if
11:end for
4.2 Continuous Extension of the PCBF
As the PCBF $h_{\mathrm{PB}}(x)=0$ for all $x\in\mathcal{S}_{\mathrm{PB}}$, errors in its approximation can lead to undesired conservativeness within $\mathcal{S}_{\mathrm{PB}}$ if $\hat{h}_{\mathrm{PB}}(x)>0$, as potentially safe control inputs can be deemed unsafe. In order to avoid such unnecessary safety filter interventions, we propose a continuous extension of $h_{\mathrm{PB}}$ in $\mathcal{S}_{\mathrm{PB}}$, which is negative. By sampling this continuous extension within $\mathcal{S}_{\mathrm{PB}}$ rather than $h_{\mathrm{PB}}$, the data used for training is guaranteed to be negative rather than 0 and therefore mitigates potential learning errors leading to states being considered unsafe. Consider the function
$h_{\mathrm{CE}}(x):=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}},\\ \bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}},\end{cases}$ (13)
with $h_{\mathrm{CE}}:\mathcal{D}_{\mathrm{PB}}\rightarrow\mathbb{R}$ and where we define
$\bar{h}_{\mathrm{PB}}(x(k)):=\min_{x_{i|k},u_{i|k},\xi_{i|k}}\;\;$ $\max_{i,j}\{[\xi_{i|k}]_{j},\xi_{N|k}\}$ (14a)
s.t. $\forall\,i=0,\dots,N-1,$
$x_{0|k}=x(k),$ (14b)
$x_{i+1|k}=f(x_{i|k},u_{i|k},0),$ (14c)
$u_{i|k}\in\mathcal{U},$ (14d)
$x_{i|k}\in\overline{\mathcal{X}}_{i}(\xi_{i|k}),$ (14e)
$h_{f}(x_{N|k})\leq\xi_{N|k}.$ (14f)
Note that the slacks in (14e) and (14f) are not required to be positive, implying negativity of $\bar{h}_{\mathrm{PB}}$ when the state and terminal constraints are satisfied. This formulation can be solved using an epigraph transformation and shows close resemblance to discrete-time Hamilton-Jacobi value functions such as [50]. We show that the function $h_{\mathrm{CE}}$ is continuous on $\mathcal{D}_{\mathrm{PB}}$ and a CBF according to Definition 1.
Theorem 3
It holds that $h_{\mathrm{CE}}$ as defined in (13) is a CBF according to Definition 1.
Proof 4.4.
We first prove continuity of $h_{\mathrm{CE}}$ through continuity of $h_{\mathrm{PB}}$ on $\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}\mathcal{S}_{\mathrm{PB}}$, continuity of $\bar{h}_{\mathrm{PB}}$ on $\mathcal{S}_{\mathrm{PB}}$ and that $h_{\mathrm{PB}}(x)=\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial\mathcal{S}_{\mathrm{PB}}$.
1) Continuity of $h_{\mathrm{PB}}$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}\mathcal{S}_{\mathrm{PB}}$: In [12, Theorem III.6 b)], it is shown that $h_{\mathrm{PB}}$ is continuous for all $x\in\mathcal{D}_{\mathrm{PB}}$, which implies continuity of $h$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$.
2) Continuity of $\bar{h}_{\mathrm{PB}}$ for all $x\in\mathcal{S}_{\mathrm{PB}}$: We can show continuity of $\bar{h}_{\mathrm{PB}}$ by using similar arguments as in [12, Theorem III.6 b)]. We show that for any $\epsilon>0$, there exists $\delta>0$ such that $\|x-\bar{x}\|<\delta$ implies $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$. Consider the optimal solution of (14) at $x$, i.e. $u_{i|k}^{*}(x)$ as a candidate solution in $\bar{x}$, i.e. $\bar{u}_{i|k}(\bar{x})=u^{*}_{i|k}$, which results in the state predictions $\bar{x}_{i|k}(\bar{x})$ with $\bar{x}_{0|k}=\bar{x}$. The resulting slacks are then given by $\bar{\xi}_{i|k}(\bar{x})=\min\{0,c(\bar{x}_{i|k})+\Delta_{i}\mathbf{1}\}$ and $\bar{\xi}_{N|k}=\min\{0,h_{f}(\bar{x}_{N|k}(\bar{x}))\}$. As the initial state and the corresponding inputs lie in a compact set, i.e. $\bar{x}\in\mathcal{S}_{\mathrm{PB}}$ and $\bar{u}_{i|k}(\bar{x})\in\mathcal{U}$, and the functions $f,c_{j}$ and $h_{f}$ are continuous, it holds that the candidate solution also lies within a compact set from [12, Lemma C.3] and from the Heine-Cantor theorem, it follows that $f,c_{j}$ and $h_{f}$ are uniformly continuous. As compositions of functions, the minimum and additions of uniform continuous functions all preserve uniform continuity, the state predictions, the candidate slack variables and the corresponding constructed candidate value function $\tilde{h}_{\mathrm{PB}}(\bar{x})=\max_{i,j}\{[\bar{\xi}_{i|k}(\bar{x})]_{j},\bar{\xi}_{N|k}(\bar{x})\}$ are uniformly continuous in the initial condition $\bar{x}$. From the fact that the optimal solution $\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(\bar{x})$ due to the minimisation in (14), it holds that for any $\epsilon>0$, there exists a uniform $\delta>0$ with $\|x-\bar{x}\|<\delta$, such that $\bar{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)\leq\tilde{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)<\epsilon$. The same arguments hold when using the optimal solution at $\bar{x}$ in $x$, i.e. $\bar{u}_{i|k}(x)=u_{i|k}^{*}(\bar{x})$, such that there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})<\epsilon$, which in turn implies that for any $\epsilon>0$, there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$.
[…]
As the function $h_{\mathrm{CE}}$ is continuous on its subdomains $\mathcal{S}_{\mathrm{PB}}$ and $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and $\bar{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)=0$ for all $x\in\partial\mathcal{S}_{\mathrm{PB}}$, it holds that $h_{\mathrm{CE}}$ is continuous on $\mathcal{D}_{\mathrm{PB}}$. As $h(x)=h_{\mathrm{PB}}(x)$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and $h_{\mathrm{CE}}(x)\leq 0$ for all $x\in\mathcal{S}_{\mathrm{PB}}$, the conditions in Definition 1 are satisfied.
Remark 4.5.
Note that the function
[…]
In principle, this continuous extension requires solving two optimisation problems for each sample $x\in\mathcal{S}_{\mathrm{PB}}$, i.e., (9) to determine if $x\in\mathcal{S}_{\mathrm{PB}}$ or $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and (14) in case $x\in\mathcal{S}_{\mathrm{PB}}$. However, a geometric sampler such as used in [41] can be used, which performs a line search over given seed and target points. By using such a geometric sampler, information of the previous solution can be used to solve either (9) or (14) based on which problem was previously solved. It follows that two optimisation problems only need to be solved if the samples on the line cross over from $\mathcal{S}_{\mathrm{PB}}$ into $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ or vice-versa. Additionally, this geometric sampler benefits from being able to warmstart the solver using the previous optimal solution $\mathbf{o}=\{x^{*}_{i},u^{*}_{i},\xi^{*}_{i},\bm{\nu}^{*},\bm{\lambda}^{*}\}$, where $\bm{\nu}^{*}$ and $\bm{\lambda}^{*}$ denote the optimal dual variables. Given that the optimal solution is shown to be continuous in $x$ in [12, Theorem III.6] and Theorem 3, the previously computed optimal solution will be close to the solution of the next iteration in the line search. This allows the computation time during sampling to be significantly reduced as shown in [41]. Finally, the random walk can be parallelised by performing line search for different goal points given a random starting seed. The resulting sampling algorithm is then given by Algorithm 2 in the Appendix.